# Light of Normals: Unified Feature Representation for Universal Photometric Stereo

### v4

@@ Abstract
Figure 1: (left) Given multi-light images from a fixed viewpoint, LINO UniPS recovers sharper, more faithful normals than UniPS/SDM-UniPS and visually rivals a 3D scanner. (right) On the DiLiGenT, a clear correlation exists between the consistency of encoder features (CSIM/SSIM) and the final reconstruction accuracy (1/MAE).
@@ 1 Introduction
The first challenge lies in the ineffective decoupling of illumination and normal cues. Existing encoders jointly process lighting and normal, but without explicit illumination representation the decoder inherits unstable features, producing inconsistent normal predictions. Interestingly, we observe that when encoders produce more similar lighting features across different inputs, the accuracy of normal estimation improves (Fig. 1 right), suggesting that explicitly factoring out illumination is crucial for normal recovery. The second challenge is the loss of geometry details. Conventional up/downsampling operations (e.g., bilinear interpolation Ikehata (2022), pixel shuffle Ikehata (2023)) either smooth fine details or disrupt high-frequency semantics Odena et al. (2016), leading to degraded normal quality, especially in regions with complex variation (see Fig. 1 left).
In this paper, we propose LINO UniPS, a ViT-based framework that explicitly decouples illumination from surface cues while preserving geometric detail. First, we introduce learnable Light Register Tokens to capture illumination information. To handle diverse lights, we design three distinct types of tokens, each dedicated to a specific light source: point, direction light, and environment (Env) lights. Our key innovation is an explicit light alignment supervision, which guides each register token to learn the characteristics of its corresponding illumination type. Second, we introduce an Interleaved Attention Block to process the image features alongside these specialized registers. This block employs a global cross-image attention mechanism that simultaneously attends to all tokens from all lighting conditions. The synergy between the Light Register Tokens and the Interleaved Attention Block enables the decoupling of lighting and normal cues, allowing our encoder to yield a unified feature representation. In parallel, we introduce two components to ensure the preservation of fine-scale geometry: a Wavelet-based Dual-branch Architecture, which introduces the wavelet domain to help the model better attend to high-frequency details, and a Normal-gradient Perception Loss that more heavily penalizes errors in high-frequency regions during training, thereby further refining local predictions. We further construct PS-Verse, a new dataset graded by surface complexity and lighting diversity, which enhances robustness and generalization under challenging real-world lighting. Together, these designs enable state-of-the-art accuracy, robustness, and generalization across synthetic and real benchmarks.
@@ 2 Related Work
Calibrated Photometric Stereo: Calibrated PS Woodham (1980) deduces surface normals by assuming meticulously pre-calibrated illumination, often relying on precise light source parameters. This stringent requirement limits the applicability of early methods. Initial methods successfully addressed Lambertian surfaces under such known lighting Woodham (1980). Subsequent research extended capabilities to handle complex non-Lambertian reflectance, including varied BRDFs and specular highlights Grossberg & Nayar (2003); Chandraker et al. (2005); Goldman et al. (2005; 2010); Ikehata et al. (2012), albeit always presupposing perfectly known illumination conditions. More recently, deep learning techniques Jung et al. (2015); Santo et al. (2017); Chen et al. (2018); Ju et al. (2021) have significantly enhanced the ability to model intricate materials within this calibrated framework. However, the reliance of Calibrated PS on precise light source parameters severely restricts its applications to controlled laboratory environments, thereby limiting its real-world utility.
Uncalibrated Photometric Stereo: To alleviate photometric stereo’s dependence on pre-calibrated light sources, the task of Uncalibrated Photometric Stereo (Uncalibrated PS) was introduced. Uncalibrated PS methods Hayakawa (1994); Belhumeur & Kriegman (1996); Mallick et al. (2005); Chen et al. (2020); Taniai & Maehara (2018); Kaya et al. (2021); Lichy et al. (2022) commonly adopt a two-stage pipeline: an initial stage where a network module estimates the unknown lighting parameters, followed by a second stage where these estimates are utilized as known inputs, like conventional Calibrated PS frameworks. While such Uncalibrated PS techniques have demonstrated success in handling more complex scenarios Lichy et al. (2022), their application under unconstrained natural or ambient illumination encounters significant challenges. These challenges primarily stem from the inherent difficulty in accurately modeling the physics of such arbitrary and often intricate lighting environments.
@@ 3 Method
The core task of Universal PS is to accurately recover the surface normal map $N\in\mathbb{R}^{H\times W\times 3}$ of an object or scene from a set of $F$ images $\{I_{f}\}_{f=1}^{F}$, where $I_{f}\in\mathbb{R}^{H\times W\times 3}$, captured under a single viewpoint but multiple unknown lighting conditions Ikehata (2022).
As illustrated in Fig. 2, our method processes multi-light images through an encoder-decoder architecture. Within the encoder, the pipeline consists of three main stages: 1) First, the input images undergo processing by the Wavelet Feature Extractor. This module simultaneously generates downsample components and low- and high-frequency wavelet components via wavelet decomposition Daubechies (1990); Finder et al. (2024), and then converts both components into independent token sequences through the ViT-based backbone Dosovitskiy et al. (2021). 2) These sequences then enter the Light Registered Attention Module, where they are prepended with specialized Light Register Tokens (Point, Direction, Env) to aggregate global illumination, before being processed by cascaded Interleaved Attention Blocks. 3) Next, these tokens, excluding the Light Register Tokens, are fed into the Wavelet Aggregator. Here, features from the downsample and wavelet branches are fused to produce the unified feature representation, which is finally passed to a decoder for normal prediction.
The training is guided by two supervisory signals. First, a light alignment loss ensures that the register tokens capture the characteristics of their corresponding illumination sources; this loss is assigned a small weight, allowing it to serve as a regularizer. Second, a Normal Gradient Perception Loss enhances the model’s ability to reconstruct fine-grained geometric details. Furthermore, this section only provides a high-level overview; please refer to appendix A1.1 for details.
Figure 2: Overview of the LINO UniPS architecture. The encoder includes a Wavelet Feature Extractor, a Light Registered Attention Module and a Wavelet Aggregator, which together encode and fuse wavelet and downsample domain features for obtaining unified feature. The decoder is similar to SDM UniPS Ikehata (2023).
@@ 3.1 LINO-UniPS
Drawing on the feature alignment methods used in VAVAE Yao et al. (2025) and REPA Yu et al. (2025) to accelerate generative model training, we encode the three types of Light Register Tokens to align with information about the light sources in the training dataset, including the point, direction and environment lights, (see Fig. 2), and supervise each of them separately with a cosine similarity loss. We summarize the form of the loss functions as follows:
$\mathcal{L}_{\text{light}}=\lambda_{1}\mathcal{L}_{\text{env}}+\lambda_{2}\mathcal{L}_{\text{point}}+\lambda_{3}\mathcal{L}_{\text{direction}}$
@@
Through light alignment supervision, these tokens learn to capture and encode their respective illumination information, thereby achieving decoupling of lighting from normal features. This is demonstrated by the phenomena observed in Fig. 3: the register tokens show specialized focus: Direction and Env register tokens attend to broad regions, while the Point register tokens focuse more sharply. This aligns with the expected low- and high-frequency characteristics of these light types, respectively, confirming our registers’ effectiveness.
Interleaved Attention Block: Previous methods like UniPS and SDM UniPS rely solely on Frame Attention (Frame) Dosovitskiy et al. (2021) and Light Axis Attention (Light) Ikehata (2023). While these mechanisms facilitate local information flow, they are insufficient for effectively decoupling lighting variations and normal features. Inspired by VGGT Wang et al. (2025), our Interleaved Attention block introduces a more powerful global cross-image attention mechanism (Global) that attends to all input image tokens simultaneously. Specifically, our block applies four attention layers in an interleaved sequence: Frame $\rightarrow$ Light $\rightarrow$ Global $\rightarrow$ Light (see Fig.2).
The key advantage of this design is its ability to aggregate and fuse information across multiple hierarchical levels. Light-axis attention operates at the patch level, frame attention captures inter-image context, and our global attention integrates information at the cross-image level. This multi-level awareness enables the model to build a holistic understanding of the global illumination from local to global scales, thereby achieving a better decoupling from the intrinsic normal features.
@@
Figure 3: The attention maps for our different Light Register Tokens, derived from the encoder’s final-layer feature maps for both real (left) and synthetic (right) data.
@@
$\mathcal{L}_{n}=\lambda_{4}\sum(N-\tilde{N})^{2}\odot C+\lambda_{5}\sum(\tilde{G}-G)^{2},$ (1)
@@ 3.2 PS-Verse Dataset
Leveraging publicly available large-scale 3D asset datasets Deitke et al. (2022; 2023); Vecchio & Deschaintre (2024), we propose the PS-Verse dataset (see Tab. 1 for a detailed comparison with other Photometric Stereo datasets). To increase the lighting complexity in rendered scenes, we carefully select 17,805 textured 3D models with UV coordinates and PBR materials from the Objaverse dataset Deitke et al. (2022), which contains nearly 800,000 models. Following Dora Chen et al. (2024), which quantifies geometric complexity based on the density of sharp edges, we categorize the objects into four complexity levels. We then construct scenes by recursively selecting 4 to 6 objects from each level, and classify scenes into four corresponding complexity tiers.
Regarding materials, textures are randomly sampled from the MatSynth large-scale PBR material database Vecchio & Deschaintre (2024), consisting of 806 metallic, 1,226 specular, and 3,321 diffuse material groups. To balance the distribution of specular and diffuse lighting effects, we assign object materials according to the ratio of 1:4:2.5:2.5 across original object textures, diffuse, specular, and metallic materials. To more realistically simulate fine surface detail interactions with light, we introduce normal maps for objects for the first time in such data, defining this as a fifth complexity level that injects rich high-frequency lighting details.
In terms of lighting setup, we use uniform environment light to simulate non-dark ambient conditions. Detailed lighting configurations are provided in the appendix A2. Overall, the PS-Verse dataset consists of 100,000 scenes generated with Blender Foundation . Each scene includes two rendering outputs: with and without normal maps. We render 20 images at a resolution of 512 per scene.
@@ 4 Experiments
$\mathcal{L}=\mathcal{L}_{\text{light}}+\mathcal{L}_{\text{n}}$
@@
Evaluation Metric: We evaluate our results using three metrics. For surface normal accuracy, we measure the Mean Angular Error (MAE, in degrees, ↓). Then, we assess feature similarity via Cosine Similarity (CSIM ↑) on normalized features and Structural Similarity Index Wang et al. (2004) (SSIM ↑) on their 3D PCA projections.
@@ 4.1 Ablation Study and Comparison Experiment
Ablation: The main ablation results are presented in Tab. 2. First, to demonstrate the superiority of our PS-Verse dataset, we retrain the Uni MS-PS Hardy et al. (2024) on both the PS-Mix Ikehata (2023) and our PS-Verse. We then compare these models against the officially released Uni MS-PS, which was trained on its native data (PS-Uni MS-PS). The results show that the model trained on PS-Verse yields substantially better feature similarity scores (CSIM and SSIM) and a lower MAE for normal reconstruction, which strongly indicates that PS-Verse is more effective for training high-performance models. Second, we conduct ablations of the various modules within our LINO UniPS method, starting from a baseline model where all our proposed enhancements are removed. Our initial finding is that by merely incorporating unsupervised Light Register Tokens, both the feature similarity metrics and the normal reconstruction results show an improvement. This suggests that these additional tokens can often capture global lighting information even without direct supervision, thereby aiding the disentanglement of normals from illumination. Progressively incorporating our global cross-image attention mechanism to our interleaved Attention Block and light alignment supervision leads to more significant improvements in both feature similarity and normal reconstruction performance. This further validates our central conclusion: that more effectively decoupling lighting and normal features enhances the quality of normal reconstruction. Next, we integrate the wavelet branch to form our Wavelet-based Dual-branch Architecture and the Normal-gradient Perception Loss for extracting fine-grained context. While these two modules have a modest impact on feature similarity metrics, they substantially improve performance on data with complex geometries as intended. Finally, from the last row, adopting curriculum learning boosts all metrics, improving feature similarity while also reducing MAE, which confirms the effectiveness of this training strategy.
@@
Table 4: Evaluation on DiLiGenTShi et al. (2018). Uses all 96 images unless otherwise noted (K).
Method Ball Bear Buddha Cat Cow Goblet Harvest Pot1 Pot2 Reading Avg.
@@
Table 5: Evaluation on LUCES Mecca et al. (2021). Uses all 52 images unless otherwise noted (K).
Method Ball Bell Bowl Buddha Bunny Cup Die Hippo House Jar Owl Queen Squirrel Tool Avg.

### v5

@@ Abstract
Figure 1: (left) Given multi-light images from a fixed viewpoint, LINO UniPS recovers sharper, more faithful normals than UniPS/SDM UniPS and visually rivals a 3D scanner. (right) On the DiLiGenT, a clear correlation exists between the consistency of encoder features (CSIM/SSIM) and the final reconstruction accuracy (1/MAE).
@@ 1 Introduction
The first challenge lies in the ineffective decoupling of illumination and normal cues. Existing encoders jointly process lighting and normal, but without explicit illumination representation the decoder inherits unstable features, producing inconsistent normal predictions. Interestingly, we observe that when encoders produce more similar normal features across different inputs, the accuracy of normal estimation improves (Fig. 1 right), suggesting that explicitly factoring out illumination is crucial for normal recovery. The second challenge is the loss of geometry details. Conventional up/downsampling operations (e.g., bilinear interpolation Ikehata (2022), pixel shuffle Ikehata (2023)) either smooth fine details or disrupt high-frequency semantics Odena et al. (2016), leading to degraded normal quality, especially in regions with complex variation (see Fig. 1 left).
In this paper, we propose LINO UniPS, a ViT-based framework that explicitly decouples illumination from surface cues while preserving geometric detail. First, we introduce learnable Light Register Tokens to capture illumination information. To handle diverse lights, we design three distinct types of tokens, each dedicated to a specific light source: point, directional lights, and environment (Env) lights. Our key innovation is an explicit Light Alignment supervision, which guides each register token to learn the characteristics of its corresponding illumination type. Second, we introduce an Interleaved Attention Block to process the image features alongside these specialized registers. This block employs a global cross-image attention mechanism that simultaneously attends to all tokens from all lighting conditions. The synergy between the Light Register Tokens and the Interleaved Attention Block enables the decoupling of lighting and normal cues, allowing our encoder to yield a unified feature representation. In parallel, we introduce two components to ensure the preservation of fine-scale geometry: a Wavelet-based Dual-branch Architecture, which introduces the wavelet domain to help the model better attend to high-frequency details, and a Normal-gradient Perception Loss that more heavily penalizes errors in high-frequency regions during training, thereby further refining local predictions. We further construct PS-Verse, a new dataset graded by surface complexity and lighting diversity, which enhances robustness and generalization under challenging real-world lighting. Together, these designs enable state-of-the-art accuracy, robustness, and generalization across synthetic and real benchmarks.
@@ 2 Related Work
Calibrated Photometric Stereo: Calibrated PS Woodham (1980) deduces surface normals by assuming meticulously pre-calibrated illumination, often relying on precise light source parameters. This stringent requirement limits the applicability of early methods. Initial methods successfully addressed Lambertian surfaces under such known lighting Woodham (1980). Subsequent research extended capabilities to handle complex non-Lambertian reflectance, including varied BRDFs and specular highlights Grossberg and Nayar (2003); Chandraker et al. (2005); Goldman et al. (2005; 2010); Ikehata et al. (2012), albeit always presupposing perfectly known illumination conditions. More recently, deep learning techniques Jung et al. (2015); Santo et al. (2017); Chen et al. (2018); Ju et al. (2021) have significantly enhanced the ability to model intricate materials within this calibrated framework. However, the reliance of Calibrated PS on precise light source parameters severely restricts its applications to controlled laboratory environments, thereby limiting its real-world utility.
Uncalibrated Photometric Stereo: To alleviate photometric stereo’s dependence on pre-calibrated light sources, the task of Uncalibrated Photometric Stereo (Uncalibrated PS) was introduced. Uncalibrated PS methods Hayakawa (1994); Belhumeur and Kriegman (1996); Mallick et al. (2005); Chen et al. (2020); Taniai and Maehara (2018); Kaya et al. (2021); Lichy et al. (2022) commonly adopt a two-stage pipeline: an initial stage where a network module estimates the unknown lighting parameters, followed by a second stage where these estimates are utilized as known inputs, like conventional Calibrated PS frameworks. While such Uncalibrated PS techniques have demonstrated success in handling more complex scenarios Lichy et al. (2022), their application under unconstrained natural or ambient illumination encounters significant challenges. These challenges primarily stem from the inherent difficulty in accurately modeling the physics of such arbitrary and often intricate lighting environments.
@@ 3 Method
The task of Universal PS is to accurately recover the surface normal map $N\in\mathbb{R}^{H\times W\times 3}$ of an object or scene from a set of $F$ images $\{I_{f}\}_{f=1}^{F}$, where $I_{f}\in\mathbb{R}^{H\times W\times 3}$, captured under a single viewpoint but multiple unknown lighting conditions $\{\mathbf{L}_{f}\}_{f=1}^{F}$ Ikehata (2022).
As illustrated in Fig. 2, our method processes multi-light images through an encoder-decoder architecture. Within the encoder, the pipeline consists of three main stages: 1) First, the input images undergo processing by the (I) Wavelet Feature Extractor. This module simultaneously generates downsample components and low- and high-frequency wavelet components via wavelet decomposition Daubechies (1990); Finder et al. (2024), and then converts both components into independent token sequences through the backbone Dosovitskiy et al. (2021). 2) These sequences then enter the (II) Light Registered Attention Module, where they are prepended with specialized Light Register Tokens (Point, Direction, Env) to aggregate global illumination, before being processed by cascaded Interleaved Attention Blocks. 3) Next, these tokens, excluding the Light Register Tokens, are fed into the (III) Wavelet Aggregator. Here, features from the downsample and wavelet branches are fused to produce the unified feature representation, which is finally passed to a decoder for normal prediction.
The training is guided by two supervisory signals. First, a Light Alignment loss ensures that the register tokens capture the characteristics of their corresponding illumination sources. Second, a Normal Gradient Perception Loss enhances the model’s ability to reconstruct fine-grained geometric details. Furthermore, this section only provides a high-level overview; please refer to appendix A1.1 for details.
Figure 2: Overview of the LINO UniPS architecture. The encoder includes (I) Wavelet Feature Extractor, (II) Light Registered Attention Module and (III) Wavelet Aggregator, which together fuse wavelet and downsample domain features for obtaining unified feature.
@@ 3.1 LINO-UniPS
Drawing on the feature alignment methods used in VAVAE Yao et al. (2025) and REPA Yu et al. (2025) to accelerate generative model training, we encode the three types of Light Register Tokens to align with information about the light sources in the training dataset, including the point, direction and environment lights, (see Fig. 2, upper right corner), and supervise each of them separately with a cosine similarity loss. We summarize the form of the loss functions as follows:
$\mathcal{L}_{\text{light}}=\lambda_{1}\mathcal{L}_{\text{env}}+\lambda_{2}\mathcal{L}_{\text{point}}+\lambda_{3}\mathcal{L}_{\text{direction}}$ (1)
@@
Through Light Alignment supervision, these tokens learn to capture and encode their respective illumination information, thereby achieving decoupling of lighting from normal features. This is demonstrated by the phenomena observed in Fig. 3. The attention maps show that the register tokens have learned specialized focus. Specifically, the Point register tokens exhibit a sparse and sharp attentional focus, concentrating on high-intensity regions, which strongly aligns with the high-frequency characteristics of point light sources that produce distinct specular highlights. The Direction and Env register tokens attend to broad, spatially-diffuse regions across the object, aligning perfectly with the expected low-frequency characteristics of directional and environmental illumination that govern global shading and softer shadows. This clear division of labor confirms the effectiveness of our register tokens in disentangling the different physical components of the illumination.
Interleaved Attention Block: Previous methods like UniPS and SDM UniPS rely solely on Frame Attention (Frame) Dosovitskiy et al. (2021) and Light Axis Attention (Light) Ikehata (2023). While these mechanisms facilitate local information flow, they are insufficient for effectively decoupling lighting variations and normal features. Inspired by VGGT Wang et al. (2025), our Interleaved Attention block introduces a more powerful global cross-image attention mechanism (Global) that attends to all input image tokens simultaneously. Specifically, our block applies four attention layers in an interleaved sequence: Frame $\rightarrow$ Light $\rightarrow$ Global $\rightarrow$ Light (see Fig. 2, middle).
The key advantage of this design is its ability to aggregate and fuse information across multiple hierarchical levels. Light-axis attention operates at the patch level, frame attention captures intra-image context, and our global attention integrates information at the cross-image level. This multi-level awareness enables the model to build a holistic understanding of the global illumination from local to global scales, thereby achieving a better decoupling from the intrinsic normal features.
@@
Figure 3: Attention maps for our different Light Register Tokens, derived from the encoder’s final-layer feature maps for both real (left) and synthetic (right) data. These maps visually confirm that the tokens learn to focus on distinct distributions of illumination information.
@@
$\mathcal{L}_{\text{n}}=\lambda_{4}\sum(N-\tilde{N})^{2}\odot C+\lambda_{5}\sum(\tilde{G}-G)^{2},$ (2)
@@ 3.2 PS-Verse Dataset
Leveraging publicly available large-scale 3D asset datasets Deitke et al. (2022; 2023); Vecchio and Deschaintre (2024), we propose the PS-Verse dataset (see Tab. 1 for a detailed comparison with other Photometric Stereo datasets). To increase the lighting complexity in rendered scenes, we carefully select 17,805 textured 3D models with UV coordinates and PBR materials from the Objaverse dataset Deitke et al. (2022), which contains nearly 800,000 models. Following Dora Chen et al. (2024), which quantifies geometric complexity based on the density of sharp edges, we categorize the objects into four complexity levels. We then construct scenes by recursively selecting 4 to 6 objects from each level, and classify scenes into four corresponding complexity tiers.
Regarding materials, textures are randomly sampled from the MatSynth large-scale PBR material database Vecchio and Deschaintre (2024), consisting of 806 metallic, 1,226 specular, and 3,321 diffuse material groups. To balance the distribution of specular and diffuse lighting effects, we assign object materials according to the ratio of 1:4:2.5:2.5 across original object textures, diffuse, specular, and metallic materials. To more realistically simulate fine surface detail interactions with light, we introduce normal mapping for objects for the first time in such data, defining this as a fifth complexity level that injects rich high-frequency lighting details.
In terms of lighting setup, we use uniform environment light to simulate non-dark ambient conditions. Detailed lighting configurations are provided in the appendix A1.3.3. Overall, the PS-Verse dataset consists of 100,000 scenes generated with Blender Foundation . Each scene includes two rendering outputs: with and without normal mapping. We render 20 images at a resolution of 512 per scene.
@@ 4 Experiments
$\mathcal{L}=\mathcal{L}_{\text{light}}+\mathcal{L}_{\text{n}}$ (3)
@@
Figure 6: Qualitative results for the main ablation study. The visual comparisons are shown on samples from our PS-Verse Testdata. We recommend zooming in to observe fine-grained details.
Evaluation Metric: We evaluate our results using three metrics. For surface normal accuracy, we measure the Mean Angular Error (MAE, in degrees, ↓). To assess feature similarity, we utilize Cosine Similarity (CSIM ↑) on normalized features and the Structural Similarity (SSIM ↑) Wang et al. (2004) on their 3D PCA projections. For both metrics, the final score is obtained by averaging the similarity values computed between all possible feature pairs. For instance, given a set of 6 features, we calculate all $C_{6}^{2}$ pairwise values to derive this mean.
@@ 4.1 Ablation Study and Comparison Experiment
Ablation: The main ablation quantitative results are presented in Tab. 2, and the qualitative results are in Fig. 6. First, to demonstrate the superiority of our PS-Verse dataset, we retrain the Uni MS-PS Hardy et al. (2024) on both the PS-Mix Ikehata (2023) and our PS-Verse. We then compare these models against the officially released Uni MS-PS, which was trained on its native data (PS-Uni MS-PS). The results show that the model trained on PS-Verse yields substantially better feature similarity scores (CSIM and SSIM) and a lower MAE for normal reconstruction, which strongly indicates that PS-Verse is more effective for training high-performance models.
@@
Second, we conduct ablations of the various modules within our LINO UniPS method, starting from a baseline model where all our proposed enhancements are removed. Our initial finding is that by merely incorporating unsupervised Light Register Tokens, both the feature similarity metrics and the normal reconstruction results show an improvement. This suggests that these additional tokens can often capture global lighting information even without direct supervision, thereby aiding the disentanglement of normals from illumination. Progressively incorporating our global cross-image attention mechanism to our interleaved Attention Block and light alignment supervision leads to more significant improvements in both feature similarity and normal reconstruction performance. This further validates our central conclusion: that more effectively decoupling lighting and normal features enhances the quality of normal reconstruction. Next, we integrate the wavelet branch to form our Wavelet-based Dual-branch Architecture and the Normal-gradient Perception Loss for extracting fine-grained context. While these two modules have a modest impact on feature similarity metrics, they substantially improve performance on data with complex geometries as intended. Finally, from the last row, adopting curriculum learning boosts all metrics, improving feature similarity while also reducing MAE, which confirms the effectiveness of this training strategy.
@@
Table 4: Evaluation on DiLiGenT Shi et al. (2018). Uses all 96 images unless otherwise noted (K). Normal reconstruction accuracy is evaluated by Mean Angular Error (MAE).
Method Ball Bear Buddha Cat Cow Goblet Harvest Pot1 Pot2 Reading Avg. MAE
@@
Table 5: Evaluation on Luces Mecca et al. (2021). Uses all 52 images unless otherwise noted (K). Normal reconstruction accuracy is evaluated by Mean Angular Error (MAE).
Method Ball Bell Bowl Buddha Bunny Cup Die Hippo House Jar Owl Queen Squirrel Tool Avg. MAE